Suppose that the money supply in Atlantis is given by Md= 50,000-500,000(1/p) and that there are $10,000 of currency currently in circulation. (amount of currency printed by the central bank, NOT THE TOTAL MONEY SUPPLY). If the price level is given by p=$50, what is the reserve ratio in Atlantis.
Answer #1

Md= 50,000-500,000(1/p) (Money supply in Atlantis)

Price=$50

then, Md= 50,000-500,000(1/50)

   =50,000-10,000

   =$40,000

This implies that the total money supply in Atlantis is $40,000

While the currency in circulation is $10000 that means that reserve taken out by the commercial banks would be = $40,000-$10,000= $30,000

Reserve ratio= 30,000/40,000

   =0.75

This means that 0.75 of total deposits would be taken out for reserve in Atlantis.